OpenStudy (anonymous):

teesha is in the french club. there are 29 students in the club. the french teacher will pick 2 students at random to guide visiting students from france. what is the probability that teesha will not be picked as a guide?

OpenStudy (mertsj):

\[\frac{28}{29}\times\frac{27}{28}\]